OpenStudy (haaileyyy):

I think I have the right answer but if i get another question wrong I'm going to fail the test so i want to check my answer. I think the answer is -2ab-a-2 (-4ab-a-7)-(-2ab-5)

OpenStudy (rolldextro):

a (-2 b-1)-2

OpenStudy (mariamikayla):

(-4ab-a-7)-(-2ab-5)= -4ab-a-7+2ab+5= = -2ab-a-2= = -(2ab+a+2)

OpenStudy (rolldextro):

Its the same thing. -2ab-a-2=-2ab-a-2: just distribute the negative sign.